OpenStudy (anonymous):

Compute the relative rate of diffusion of ethanol (C2H5OH) to propane (C3H8).

OpenStudy (aaronq):

Graham's Law \(\huge \sf \dfrac{rate_1}{rate_2}=\sqrt{\dfrac{M_2}{M_1}}\) where M is molar mass